A solid cylinder of radius R=5.00mm carries a uniform current density of 1.50*10^6 A/m2. What is the magnitude of the magnetic field at a distance R/2 from the center axis of the cylinder? The answer is 2.36mT but can you show me the work involved? Thanks
Answer #1

r = distance from the center = R/2 = 5/2 = 2.5 mm = 2.5 x 10-3 m

= Current density = 1.50 x 106 A/m2

Using ampere's circuital law

B (2r) = (r2)

B (2) = (r)

B (2) = (12.56 x 10-7) (2.5 x 10-3) (1.50 x 106)

B = 2.36 x 10-3 T

B = 2.36 mT